### v1

# Parametric Graph Representations in the Era of Foundation Models: A Survey and Position

Abstract
Graphs have been widely used in the past decades of big data and AI to model comprehensive relational data. When analyzing a graph’s statistical properties, graph laws serve as essential tools for parameterizing its structure. Identifying meaningful graph laws can significantly enhance the effectiveness of various applications, such as graph generation and link prediction. Facing the large-scale foundation model developments nowadays, the study of graph laws reveals new research potential, e.g., providing multi-modal information for graph neural representation learning and breaking the domain inconsistency of different graph data. In this survey, we first review the previous study of graph laws from multiple perspectives, i.e., macroscope and microscope of graphs, low-order and high-order graphs, static and dynamic graphs, different observation spaces, and newly proposed graph parameters. After we review various real-world applications benefiting from the guidance of graph laws, we conclude the paper with current challenges and future research directions.
Figure 1: A Toy Example of Illustrating Graph Laws in the Real World: Different types of graphs follow some shared properties that can be used to parameterize their structure. The left figure depicts a social network, and the right figure represents flight routes between several major cities after COVID-19, where flights gradually recover to their pre-pandemic state. Both graphs exhibit two key properties: (1) Global Shrinking Law: As the graph evolves, the diameter decreases, indicating a more connected structure. (2) Local Triangle Closure Law: New links are more likely to close triangles. These behaviors are indicative of typical graph dynamics seen in social networks, transport systems, and other domains of graphs. These graph laws have the potential to be quantified and serve as fundamental principles for developing graph foundation models.
1 Introduction
In the era of big data and AI, graphs are popular data structures for modeling the complex relationships between entities. Also, graph-based research (e.g., graph mining and graph representation learning) provides the foundation for many real-world applications, such as recommender system [46, 22, 49, 51], social network analysis [35, 23], information retrieval [36, 45, 32], anomaly detection [63, 64, 17], natural language processing [50, 7, 65], computer vision [8, 27], AI4Science [59, 18, 15], etc.
Figure 2: Position of Graph Law in Graph Representations.
To model those real-world tasks within graphs, graph representations are indispensable middleware that provides the basis for specific and complex task-oriented computations. To be specific, graph representations can be decomposed into three aspects, (1) graph embedding (i.e., vector representation), (2) graph law (i.e., parametric representation), and (3) graph visualization (i.e., visual representation). First of all, graph representations can be in the form of embedding matrices, i.e., the graph topological information and attributes are encoded into matrices, which has been widely discussed and studied in the research community and usually can be referred to as graph representation learning[20]. Then, graphs can be also represented by plotting directly for a better human-understandable illustration. For example, one interesting research topic is how to plot the graph topological structures into the 2D space with less structure distortion. More interesting works can be referred to [16]. Last but not least, graphs can also be represented by a few key parameters such as Erdős-Rényi random graph $G(n,p)$ [12], where $n$ stands for the number of nodes in the graph, and $p$ stands for the independent edge connection probability in the graph. As shown in Figure 2, these three representations can have overlapping to mutually contribute to each other [16].
1.1 Motivation of this Paper
In the modern graph deep learning community, graph embedding (also referred to as graph representation learning) has attracted unprecedented research attention varying from unsupervised (e.g., Node2Vec [19]) to semi-supervised (e.g., GCN [28]), and also the novel neural architecture from transformer (e.g., GraphTransformer [13]). However, the graph law (also referred to as graph parametric representation) has great research potential, yet the corresponding research stays in a not fully exploited stage.
At the beginning, we would point out “What is Graph Law or Graph Parametric Representation?”. In general, it is referred to as finding some key parameters and their relations to describe the given graph, which description is expected to preserve the entire or part of the property of the given graph. According to the previous research [33, 34, DBLP:conf/kdd/FuFMTH22], the rigorous graph law (or graph parametric representation) relies two fundamental steps. The first step is to decide which parameters we use to describe the graph, e.g., the node degree. The second step is to observe the graph in a statistical manner and then determine the value of parameters, the distribution of parameters, the relation among parameters, etc. For example, the relation between the possibility of a newly-arrived node connecting to an old node (parameter 1) and the degree of that old node (parameter 2) is studied by maximum likelihood estimation (MLE) based on the observed real-world graph data [34].
Then, we discuss why studying the graph parametric representation is important in the graph research community nowadays with the following two concrete examples.
  • Foundation of Graph Foundation Models. Similar to building the foundation models in other modalities [61, 62, 3], Graph data integration is the base to support various graph AI developments by aggregating different domain graph data for extracting non-trivial, abundant, and useful knowledge. But its realization is very challenging, for example, unlike the independent and identically distributed (IID) data, the distribution of graph data from different domains can be quite different and not easy to be jointly leveraged, e.g., the graph size, the attribute dimension, and the attribute meaning of different domain graphs place a barrier for distilling a consensus intelligence [44]. Graph law acts like a potential trigger to break the inconsistency of different domains, such that different graphs can be described under the same statistical language. With a suite of powerful graph laws, the cross-domain graph (or subgraph) representation complexities can be reduced to several shareable parameters, akin to Erdős–Rényi graphs but accounting for heterogeneity and temporality, such that the large-scale foundation model training among various graph data can become promising.
  • Bring Graphs to Foundation Models. Whether the input graph statistical property gets well preserved during the representation learning process has become an increasingly interesting research question recently, and some preliminary theoretical studies [48] will be discussed in the following part of this paper. Further, take the concrete application scenarios as examples, in the climate domain, modeling the geolocation as graphs [31], graph neural networks (GNNs) have been deployed for the weather forecasting and obtain the outperformance. Although understood by machines, a follow-up question gets raised as to whether this GNN-encoded knowledge obeys the physical law or equation of climate, i.e., does the machine understand the climate as humans do? Thus, a possible solution is to find a way to model the graph law in terms of physical rules, try to encode the law along with the neural representation learning process, and see the decision variance to verify the hypothesis. In addition to natural science, graph law also has the potential to bring human activities into the neural representation learning process and boost task performance. For example, one recent study [25] shows that adding extra local topology information into the prompt can help large language models (LLMs) achieve textual node classification tasks with high accuracy. More discussion on graph laws and LLMs like [53, 21] can be found in Section 6.6.
Figure 3: Organization of Graph Law Introduction.
1.2 Organization of this Paper
Graph law is the study of investigating the statistical properties of graphs. In this survey, we introduce the graph laws studies in the macroscopic view and microscopic view, plus multiple angles like low-order and high-order connections, static and dynamic graphs, as shown in Table 1 and Figure 3.
  • Macroscopic and Microscopic Views. The macroscopic graph laws describe the graph properties in a global view, like how the total degree (or eigenvalues) distribution of the entire graph looks like [33]; while the microscopic laws try to focus on the individual behavior and investigate their behaviors as part in the entire graph [34].
  • Low-Order and High-Order Connections. Most graph laws are based on the node-level connections (i.e., low-order connections), while some graph law investigations are based on the group activities (high-order connections), i.e., motif in [40, 57], hyperedge in [10, 29], and simplex in [6, 9].
  • Static and Dynamic Graphs. Compared with static graphs, dynamic graphs allow the graph components like topology structure and node attributes to evolve over time. Correspondingly, some graph laws study how the graph parameters change over time and their temporal relations. Note that, in some graph research, the dynamics are created by the algorithms, like adding virtual nodes to preserve graph representations [1, 37]; this kind of research is beyond the discussion of this paper, and we focus on the graph law with natural time.
Moreover, in Section 4, we introduce different observation spaces and newly proposed parameters, before the corresponding laws are discovered.
Furthermore, in Section 5, we survey different real-world applications that would benefit from the guidance of graph laws, such as graph generation, link prediction, and natural language processing.
Finally, in Section 6, we conclude the survey with current challenges and 6 future directions for graph law research.
Table 1: A summary of parameteric representations of graphs. Some laws have multiple aspects and are indexed by numbers in parentheses.
Input Law Parameter Scope Order Temporality Description
Graphs Densification Law [33] Density degree $\alpha$ Macro Low Dynamic $e(t)\propto n(t)^{\alpha},\alpha\in[1,2]$, $e(t)$ is # edges at $t$
Shrinking Law [33] Effective diameter $d$ Macro Low Dynamic $d_{t+1}<d_{t}$, $d$ decreases as network grows
Motif Differing Law(1) [40] Numbers of similar motifs $n$ Macro High Dynamic $n_{1}\neq n_{2}$ for different domains
Motif Differing Law(2) [40] Motif occurring timestamp $t$ Macro High Dynamic $t_{1}\neq t_{2}$ for different motifs
Egonet Differing Law [6] Features of Egonets $X$ Macro High Static $X_{1}\neq X_{2}$ for different domains
Simplicial Closure Law [6] Simplicial closure probability $p$ Macro High Static $p$ increases with additional edges or tie strength
Spectral Power Law(1) [14] Degree, SVD, eigen distributions Macro High Static These distributions usually follow power-law
Spectral Power Law(2) [14] Degree, SVD, eigen distributions Macro High Static If one follow power-law, usually others follow
Edge Attachment Law(1) [34] Node degree $d$, edge create $p_{e}(d)$ Micro Low Dynamic $p_{e}(d)\propto d$ for node with degree $d$
Edge Attachment Law(2) [34] Node age $a(u)$, edge create $p_{e}(d)$ Micro Low Dynamic $p_{e}(d)$ seems to be non-decreasing with $a(u)$
Triangle Closure Law(1) [24] Triangular connections $e_{1},e_{2},e_{3}$ Micro Low Dynamic Strong $e_{3}\Rightarrow$ unlikely $e_{1}/e_{2}$ will be weakened
Triangle Closure Law(2) [24] Triangular connections $e_{1},e_{2},e_{3}$ Micro Low Dynamic Strong $e_{1}/e_{2}\Rightarrow$ unlikely they will be weakened
Local Closure Law [54] Local closure coefficient $H(u)$ Micro Low Static Please refer to Section 4 for details
Spectral Density Law [11] Density of states $\mu(\lambda)$ Macro High Static Please refer to Section 4 for details
Motif Activity Law(1) [57] Motif type Micro High Dynamic Motifs do not transit from one type to another
Motif Activity Law(2) [57] Motif re-appear rate Micro High Dynamic Motifs re-appear with configured rates
Hypergraphs Degree Distribution Law [10] Node degree, edge link probability Macro High Dynamic High-degree nodes are likely to form new links
SVD Distribution Law [10] Singular value distribution Macro High Static Singular value distribution usually heavy-tailed
Diminishing Overlaps [29] density of interactions $DoI(\mathcal{H}(t))$ Macro High Dynamic Overall hyperedge overlaps decrease over time
Densification Law [29] Density degree $\alpha$ Macro High Dynamic $e(t)\propto n(t)^{\alpha},\alpha\geq 1$, $e(t)$ is # hyperedges at $t$
Shrinking Law [29] Hypergraph effective diameter $d$ Macro High Dynamic $d_{t+1}<d_{t}$, $d$ decreases as network grows
Edge Interacting Law [9] Edge interacting rate Micro High Dynamic Temporally adjacent interactions highly similar
Heterographs Densification Law [47] Density degree $\alpha$, # meta-path Macro Low Dynamic $e(t)\propto n(t)^{\alpha},\alpha\geq 1$ for some meta-path
Non-densification Law [47] Density degree $\alpha$, # meta-path Macro Low Dynamic Maybe, for some meta-path, $e(t)\not\propto n(t)^{\alpha}$
2 Macroscopic Graph Laws
In this section, we introduce the graph laws from the macroscope and microscope. In detail, we will introduce what is the intuition of researchers proposing or using graph statistical properties as parameters and how they fit the value of parameters against real-world observations.
Several classical theories model the growth of graphs, for example, Barabasi-Albert model [4, 5] assumes that the graphs follow the uniform growth pattern in terms of the number of nodes, and the Bass model [38] and the Susceptible-Infected model [2] follow the Sigmoid growth (more random graph models can be founded in [12]). However, these pre-defined graph growths have been tested that they could not handle the complex real-world network growth patterns very well [30, 56]. To this end, researchers begin to fit the graph growth on real-world networks directly, to discover graph laws.
2.1 Low-Order Macroscopic Laws
Based on fitting nine real-world temporal graphs from four different domains, the authors in [33] found two temporal graph laws, called (1) Densification Laws and (2) Shrinking Diameters, respectively. First, the densification law states as follows.
$e(t)\propto n(t)^{\alpha}$ (1)
where $e(t)$ denotes the number of edges at time $t$, $n(t)$ denotes the number of nodes at time $t$, $\alpha\in[1,2]$ is an exponent representing the density degree. The second law, shrinking diameters, states that the effective diameter is decreasing as the network grows, in most cases. Here, the diameter means the node-pair shortest distance, and the effective diameter of the graph means the minimum distance $d$ such that approximately 90% of all connected pairs are reachable by a path of length at most $d$. Later, in [56], the densification law gets in-depth confirmed on four different real social networks, the research shows that the number of nodes and number of edges both grown exponentially with time, i.e., following the power-law distribution.
2.2 High-Order Macroscopic Laws
Above discoveries are based on the node-level connections (i.e., low-order connections), then several researchers start the investigation based on the group activities, for example, motifs [40], simplices [6] and hyperedges [10, 29]. Motif is defined as a subgraph induced by a sequence of selected temporal edges in [40], where the authors discovered that different domain networks have significantly different numbers of similar motifs, and different motifs usually occur at different time. Similar laws are also discovered in [6] that the authors study 19 graph data sets from domains like biology, medicine, social networks, and the web, to characterize how high-order structure emerges and differs in different domains. They discovered that the higher-order Egonet features can discriminate the domain of the graph, and the probability of simplicial closure events typically increases with additional edges or tie strength.
In hypergraphs, each hyperedge could connect an arbitrary number of nodes, rather than two [10], where the authors found that real-world static hypergraphs obey the following properties: (1) Giant Connected Components, that there is a connected component comprising a large proportion of nodes, and this proportion is significantly larger than that of the second-largest connected component. (2) Heavy-Tailed Degree Distributions, that high-degree nodes are more likely to form new links. (3) Small Effective Diameters, that most connected pairs can be reached by a small distance (4) High Clustering Coefficients, that the global average of local clustering coefficient is high. (5) Skewed Singularvalue Distributions, that the singular-value distribution is usually heavy-tailed. Later, the evolution of real-word hypergraphs is investigated in [29], and the following laws are discovered.
  • Diminishing Overlaps: The overall overlaps of hyperedges decrease over time.
  • Densification: The average degrees increase over time.
  • Shrinking Diameter: The effective diameters decrease over time.
To be specific, given a hypergraph $G(t)=(V(t),E(t))$, the density of interactions is stated as follows.
$DoI(G(t))=\frac{|~{}\{\{e_{i},e_{j}\}~{}|~{}e_{i}\cap e_{j}\neq\emptyset~{}\text{for}~{}e_{i},e_{j}\in E(t)\}~{}|}{\{\{e_{i},e_{j}\}|e_{i},e_{j}\in E(t)\}}$ (2)
and the densification is stated as follows.
$|E(t)|\propto|V(t)|^{s}$ (3)
where $s>1$ stands for the density term.
In heterogeneous information networks (where nodes and edges can have multiple types), the power law distribution is also discovered [47]. For example, for the triplet "author-paper-venue" (i.e., A-P-V), the number of authors is power law distributed w.r.t the number of A-P-V instances composed by an author.
3 Microscopic Graph Laws
In contrast to representing the whole distribution of the entire graph, many researchers try to model individual behavior and investigate how they interact with each other to see the evolution pattern microscopically.
3.1 Low-Order Microscopic Laws
In [34], the authors view temporal graphs in a three-fold process, i.e., node arrival (determining how many nodes will be added), edge initiation (how many edges will be added), and edge destination (where are each edge will be added). They ignore the deletion of nodes and edges, and they assign variables (models) to parameterize this process.
  • Edge Attachment with Locality (an inserted edge closing an open triangle): It is responsible for the edge destination.
  • Node Lifetime and Time Gap between Emitting Edges: It is responsible for edge initiation.
  • Node Arrival Rate: It is responsible for the node arrival.
To model the individual behaviors, there are many candidate models for selection. For example, in edge attachment, the probability of a newcomer $u$ to connect the node $v$ can be proportional to $v$’s current degree or $v$’s current age or the combination. Based on fitting each model to the real-world observation under the supervision of MLE principle, the authors empirically choose the random-random model for edge attachment with locality, i.e., first, let node $u$ choose a neighbor $v$ uniformly and let $v$ uniform randomly choose $u$’s neighbor $w$ to close a triangle. And node lifetime and time gap between emitting edges are defined as follows.
$a(u)=t_{d(u)}(u)-t_{1}(u)$ (4)
where $a(u)$ stand for the age of node $u$, $t_{k}(u)$ is the time when node $u$ links its $k^{th}$ edge, $d_{t}(u)$ denote the degree of node $u$ at time $t$, and $d(u)=d_{T}(u)$. $T$ is the final timestamp of the data.
$\delta_{u}(d)=t_{d+1}(u)-t_{d}(u)$ (5)
where $\delta_{u}(d)$ records the time gap between the current time and the time when that node emits its last edge. Finding the node arrival is a regression process in [34], for example, in Flickr graph $N(t)=exp(0.25t)$, and $N(t)=3900t^{2}+76000t-130000$ in LinkedIn graph.
In [52, 41], the selection of edge attachment gets flourished where the authors propose several variants of edge attachment models for preserving the graph properties. With respect to the triangle closure phenomenon, several in-depth researches follow up. For example, in [24], researchers found that (1) the stronger the third tie (the interaction frequency of the closed edge) is, the less likely the first two ties are weakened; (2) when the stronger the first two ties are, the more likely they are weakened.
3.2 High-Order Microscopic Laws
Hypergraph ego-network [9] is a structure defined to model the high-order interactions involving an individual node. The star ego-network $T(u)$ is defined as follows.
$T(u)=\{s:(u\in s)\},\forall s\in S$ (6)
where S is the set of all hyperedges (or simplices). Also, in  [9], there are other hypergraph ego-networks, like radial ego-network $R(u)$ and contracted ego-network $C(u)$. The relationship between them is as follows.
$T(u)\subseteq R(u)\subseteq C(u)$ (7)
In [9], authors observe that contiguous hyperedges (simplices) in an ego-network tend to have relatively large interactions with each other, which suggests that temporally adjacent high-order interactions have high similarity, i.e., the same nodes tend to appear in neighboring simplices.
In [57], authors try to model the temporal graph growth in terms of motif evolution activities. In brief, this paper investigates how many motifs change and what are the exact motif types in each time interval and fits the arrival rate parameter of each type of motif against the whole observed temporal graph.
4 Some New Observation Space and Newly Discovered Graph Parameters
4.1 New Different Spaces
In [14], the power law is revisited based on the eigendecomposition and singular value decomposition to provide guidance on the presence of power laws in terms of the degree distribution, singular value (of adjacency matrix) distribution, and the eigenvalue (of Laplacian matrix) distribution. The authors [14] discovered that (1) degree distribution, singular value distribution, and eigenvalue distribution follow power law distribution in many real-world networks they collected; (2) and a significant power law distribution of degrees usually indicates power law distributed singular values and power law distributed eigenvalues with a high probability.
4.2 New Parameters
Currently, if not all, most graph law research focuses on the traditional graph properties, like the number of nodes, number of edges, degrees, diameters, eigenvalues, and singular values. Here, we provide some recently proposed graph properties, although they have not yet been tested on the scale for fitting the graph law on real-world networks.
The local closure coefficient [54] is defined as the fraction of length-2 paths (wedges) emanating from the head node (of the wedge) that induce a triangle, i.e., starting from a seed node of a wedge, how many wedges are closed. According to [54], features extracted within the constraints of the local closure coefficient can improve the link prediction accuracy. The local closure efficient of node $u$ is defined as follows.
$H(u)=\frac{2T(u)}{W^{h}(u)}$
where $W^{(h)}(u)$ is the number of wedges where $u$ stands for the head of the wedge, and $T(u)$ denotes the number of triangles that contain node $u$.
The density of states (or spectral density) [11] is defined as follows.
$\mu(\lambda)=\frac{1}{N}\sum^{N}_{i=1}\delta(\lambda-\lambda_{i}),~{}\int f(\lambda)\mu(\lambda)=\text{trace}(f(H))$ (8)
where $H$ denotes any symmetric graph matrix, $\lambda_{1},\ldots,\lambda_{N}$ denote the eigenvalues of $H$ in the ascending order, $\delta$ stands for the Dirac delta function and $f$ is any analytic test function.
5 Law-Guided Research Tasks
The discovered graph laws describe the graph property, which provides guidance to many downstreaming tasks. Some examples are discussed below.
5.1 Graph Generation
If not all, in most of graph law studies [33, 56, 10, 29, 34, 41, 57], after the law (i.e., evolution pattern) is discovered, a follow-up action is to propose the corresponding graph generative model to test whether there is a realizable graph generator could generate graphs while preserving the discovered law in terms of graph properties. Also, graph generation tasks have impactful application scenarios like drug design and protein discovery [59].
For example, in [33], the Forest Fire model is proposed to preserve the macroscopic graph law while larges preserve the discovered evolution pattern.
  • First, node $v$ first chooses an ambassador (i.e., node $w$) uniformly random, and establish a link to $w$;
  • Second, node $v$ generates a random value $x$, and selects $x$ links of node $w$, where selecting in-links $r$ times less than out-links;
  • Third, node $v$ forms links to $w$’s neighbors; this step executes recursively (neighbors of neighbors) until $v$ dies out.
This proposed Forest Fire model holds the following graph properties most of time.
  • Heavy-tailed In-degrees: The highly linked nodes can easily get reached, i.e., “rich get richer”.
  • Communities: A newcomer copies neighbors of its ambassador.
  • Heavy-tailed Out-degrees: The recursive nature produces large out-degree.
  • Densification Law: A newcomer will have a lot of links near the community of its ambassador.
  • Shrinking Diameter: It may not always hold.
In [34], authors combine the microscopic edge destination model, edge initiation model, and node arrival rate together, to model the real-world temporal network’s growth. The parameters of these three models are fitted against the partial observation. i.e., $G_{\frac{T}{2}}$, which is the half of the entire evolving graph. Then they three produce the residual part of $G_{T}^{\prime}$. Finally, the generated $G_{T}^{\prime}$ is compared with the ground truth $G_{T}$, to see if the growth pattern is fully or near fully captured by these microscopic models. The procedures are stated as follows.
  • First, nodes arrive using the node arrival function obtained from $G_{\frac{T}{2}}$;
  • Second, node $u$ arrives and samples its lifetime $a$ from the age distribution of $G_{\frac{T}{2}}$;
  • Third, node $u$ adds the first edge to node $v$ with probability proportional to node $v$’s degree;
  • Fourth, node $u$ with degree $d$ samples a time gap $\delta$ from the distribution of time gap in $G_{\frac{T}{2}}$;
  • When a node wakes up, if its lifetime has not expired yet, it creates a two-hop edge using the "random-random" triangle closing model;
  • If a node’s lifetime has expired, then it stops adding edges; otherwise, it repeats from Step 4.
The generated graph $G_{T}^{\prime}$ is tested based on the comparison with the ground truth $G_{T}$, in terms of degree distribution, clustering coefficient, and diameter distribution. Taking the Flickr graph for example, the generated graph is very similar to the ground truth with aforementioned metrics [34].
5.2 Link Prediction
To learn node representation vectors for predicting links between node pairs and contributing latent applications like recommender systems, CAW-N [48] is proposed by inserting causal anonymous walks (CAWs) into the representation learning process. The CAW is a sequence of time -aware adjacent nodes, the authors claim that the extracted CAW sequence obeys the triadic closure law. To be specific, the temporal opening and closed triangles can be preserved in the extracted CAW sequence $W$. Further, to realize the inductive link prediction, CAW-N replaces the identification of each node in $W$ with the relative position information, such that the CAW sequence $W$ is transferred into anonymous $\hat{W}$. Then, the entire $\hat{W}$ is inserted into an RNN-like model and gets the embedding vector of each node, the loss function states as follows.
$enc(\hat{W})=\textbf{RNN}(\{f_{1}(I_{CAW}(w_{i}))\oplus f_{2}(t_{i-1}-t_{i})\}_{i=0,1,\ldots,|\hat{W}|})$ (9)
where $I_{CAW}(w_{i})$ is the anonymous identification of node $i$ in $\hat{W}$, $f_{1}$ is the node embedding function realized by a multi-layer perceptron, $f_{2}$ is the time kernel function for representing a discrete time by a vector, and $\otimes$ denotes the concatenation operation. The training loss comes from predicting negative (disconnected) node pairs and positive (connected) node pairs.
Also, there are some related link prediction models based on the guidance of static graph laws during the representation learning process, for example, SEAL [58] and HHNE [47].
In the SEAL framework [58], for each target link, SEAL extracts a local enclosing subgraph around it, and uses a GNN to learn general graph structure features for link prediction. The corresponding graph parameters include but are not limited to
  • Common Neighbors: Number of common neighbors of two nodes.
  • Jaccard: Jaccard similarity on the set of neighbors of two nodes.
  • Preferential Attachment: The product of the cardinal of the sets of neighbors of two nodes.
  • Katz Index: The summarization over the collection of paths of two nodes.
5.3 Natural Language Processing
To obtain the semantic representation vector of each word in the corpus, GloVe [42] is proposed, which has been considered as one of the most popular word embedding models. GloVe utilizes the power law distribution constraint during the representation learning process. $X_{ij}$ denotes the number of times that word $j$ occurs in the context of word $i$, and it follows
$X_{ij}=\frac{k}{(r_{ij})^{\alpha}}$ (10)
where $r_{ij}$ denotes the frequency rank of the word pair $i$ and $j$ in the whole corpus, and $k$ and $\alpha$ are constant hyperparameters. Then the loss function of GloVe is stated as follows
$J=\sum^{V}_{i,j}f(X_{ij})(w_{i}^{\top}w_{j}+b_{i}+b_{j}-logX_{ij})^{2}$ (11)
where $w$ is the word vector, and $b$ is the bias vector.
6 Future Directions
In this section, we would like to list several interesting research directions of graph parametric representation in modern graph research.
6.1 Graph Laws on Temporal Graphs
Discovering accurate temporal graph laws from real-world networks heavily relies on the number of networks and the size of networks (e.g., number of nodes, number of edges, and time duration). However, some of the temporal graph law studies mentioned above usually consider the number of graphs ranging from 10 to 20, when they discover the evolution pattern. The existence of time-dependent structure and feature information increases the difficulty of collecting real-world temporal graph data. To obtain robust and accurate (temporal) graph laws, we may need a considerably large amount of (temporal) network data available. Luckily, we have seen some pioneering work like TGB [26] and TUDataset [39].
6.2 Graph Laws on Heterogeneous Networks
Though many graph laws have been proposed and verified on homogeneous graphs, real-world networks are usually heterogeneous [43] and contain a large number of interacting, multi-typed components. While the existing work [47] only studied 2 datasets to propose and verify the heterogeneous graph power law, the potential exists for a transition in graph laws from homogeneous networks to heterogeneous networks, suggesting the presence of additional parameters contributing to the comprehensive information within heterogeneous networks. For example, in an academic network, the paper citation subgraph and the author collaboration subgraph may have their own subgraph laws which affect other subgraphs’ laws. Furthermore, Knowledge graphs, as a special group of heterogeneous networks, have not yet attracted much attention from the research community to study their laws.
6.3 Transferability of Graph Laws
As we can see in the front part of the paper, many nascent graph laws are described verbally without the exact mathematical expression, which hinders the transfer from the graph law to the numerical constraints for the representation learning process. One latent reason for this phenomenon is that selecting appropriate models and parameters and fitting the exact values of parameters on large evolving graphs are very computationally demanding.
6.4 Taxonomy of Graph Laws
After we discovered many graph laws, is there any taxonomy or hierarchy of those? For example, graph law A stands in the superclass of graph law B, and when we preserve graph law A during the representation, we actually have already preserved graph law B. For example, there is a hierarchy of different computer vision tasks, recently discovered [55]. And corresponding research on graph law development seems like a promising direction.
6.5 Domain-Specific Graph Laws
Since graphs serve as general data representations with extreme diversity, it is challenging to find universal graph laws that fit all graph domains because each domain may be internally different from another [60]. In fact, in many cases, we have prior knowledge about the domain of a graph, which can be a social network, a protein network, or a transportation network. Thus, it is possible to study the domain-specific graph laws that work well on only a portion of graphs and then apply the graph laws only on those graphs.
6.6 Graph Laws with LLMs
In the background of large language models (LLMs) developments, an interesting question attracts much research interest nowadays, i.e., can LLMs replace GNNs as the backbone model for graphs? To answer this question, many recent works show the great efforts [25, 53, 21], where the key point is how to represent the structural information as the input for LLMs.
For example, Instruct-GLM [53] follows the manner of instruction tuning and makes the template $\mathcal{T}$ of a 2-hop connection for a central node $v$ as follows.
[…]
where $\mathcal{A}^{v}_{k}$ represents the list of node $v$’s $k$-hop neighbors.
As discussed above, the topological information (e.g., 1-hop or 2-hop connections) can serve as external modality information to contribute to (e.g., through prompting) the reasoning ability of large language models (LLMs) [25] and achieve state-of-the-art on low-order tasks like node classification and link prediction.
Therefore, a natural question can be asked, i.e., instead of inputting local topological information to LLMs, how can we bring global topological information for LLMs to understand and make inferences for high-order tasks like graph classification, graph matching, and graph alignment? To the best of our knowledge, corresponding research still remains nascent but has great potential. Finding a proper graph parametric representation in a macroscopic way may be a viable solution for LLMs to comprehend graph-level information.
7 Conclusion
Within the survey, we first review the concepts and developing progress of graph parametric representations (i.e., graph laws) from different perspectives like microscope and microscope, low-order and high-order connections, and static and temporal graphs. We then discuss various real-world application tasks that can benefit the study of graph parametric representations. Finally, we envision the latent challenges and opportunities of graph parametric representations in modern graph research with several interesting and possible future directions.